Complaints received outside of normal business hours. Results of three delayed odor complaint response investigations concluded the stage 2 bleach pump tripped and the power outlet required maintenance. Dosing is normal since the power outlet was reset on 8/3/19. Odor scrubber trends show normal PH and normal Oxidation-Reduction Potential levels and the system is operating as intended. No abnormal odors detected during odor patrol fenceline monitoring. Typical odors detected along the east gate fenceline.